# An End-to-End Approach for Microgrid Probabilistic Forecasting and Robust Operation via Decision-focused Learning

### v1

@@ III-B1 Statistical Error Metric
III-B2 Surrogate Smart Predict-then-Optimize Loss
Traditional probabilistic forecasting metrics such as the CRPS focus solely on statistical accuracy without considering downstream decision consequences. To overcome this limitation, the traditional smart predict-then-optimize loss is introduced as a decision-focused learning paradigm that directly minimizes decision regret rather than statistical discrepancy, thereby aligning forecasting objectives with operational performance [elmachtoub2022smart]. The smart predict-then-optimize loss is shown as (10)
$)\mathcal{L}_{\text{SPO}}(\hat{D},D)=\max_{D\in\mathcal{U}_{D}}\Big(J^{*}(D)-J^{*}(\hat{D})-\nabla J^{*}(\hat{D})^{\top}(D-\hat{D}$$\Big)$ (10)
$J^{*}$ denotes the optimal value function of the operation problem, $\hat{D}$ represents the predicted load demand distribution, $D$ is the ground-truth load demand realization and $\mathcal{U}_{D}$ is the uncertainty set.
However, in practical optimization tasks, it is often non-differentiable due to the discrete and piecewise nature of the downstream optimization problem, particularly when the operation model includes binary variables or non-smooth constraints. This non-differentiability prevents gradient-based end-to-end training, hindering convergence and stability.
@@ III-B2 Surrogate Smart Predict-then-Optimize Loss
$)\mathcal{L}_{\text{SPO}}(\hat{D},D)\approx\max_{D\in\mathcal{U}_{D}}\Big(J^{*}(D)-J^{*}(\hat{D})-\nabla J^{*}(\hat{D})^{\top}(D-\hat{D}$$\Big)$ (11)
@@
Furthermore, a smooth approximation of the regret function is introduced using the softplus operator to ensure numerical stability:
@@
This smooth surrogate preserves the monotonicity of the regret term while allowing continuous gradient propagation. In the following, we denote the SPO loss function module as $\lambda\cdot\text{Regret}_{\text{SPO}}$, which represents the regret value in different prediction tasks (including load demand, PV and WT active power). Moreover, the parameter $\lambda$ is relied on to determine the weight of the regret value, which is determined by the grid search method.
This section establishes SPO as a theoretically grounded alternative to CRPS and traditional smart predict-then-optimize loss, enabling efficient gradient-based joint training of the MED and TSRO.
@@ IV Robust Operation Model
This section introduces the robust operation model used to optimize microgrid performance under uncertainty. It outlines the formulation of the optimization problem, which seeks to minimize operational costs while ensuring system resilience against fluctuations in renewable generation and load demand. The model leverages TSRO to account for uncertainty in the system’s components. Additionally, this section presents the operation constraints, including energy storage and demand-side management, and discusses the solution approach through the Column-and-Constraint Generation (C&CG) algorithm for improved computational efficiency.
@@ IV-A Objective Function
Based on the model established above, the objective function is formulated as follows,
$\min\ C_{\text{ESS}}$ $+C_{\text{WT}}+C_{\text{PV}}+C_{\text{Grid}}$
$-C_{rev}+\lambda\cdot\text{Regret}_{\text{SPO}}$ (13)
@@ IV-C1 TSRO Model
The proposed microgrid coordination model can be converted to a TSRO model, which can be formulated in the following compact matrix form:
@@
The model adopts a two-stage “min–max–min” structure, where first-stage variables $x$ represent pre-dispatch decisions (ESS and DLC plans), second-stage variables $y$ capture recourse actions, and $u$ denotes uncertain renewable and load realizations.
Therefore, we classify the three variables in the two stages as $x$, $y$ and $u$. First, $x$ represents a set of decision variables in the first stage operation, including ESS daily charging and discharging binary variables $\alpha_{\text{ch}}$}, $\alpha_{\text{dis}}$}, and DLC contracts for the entire scheduling period $K_{D,\text{con}}$}. Secondly, $y$ stands for a group of adjustable variables in the second stage including the quarter-hour ahead DLC controllable variables. $K_{\text{DLC}}$} for each quarter hour can be optimized in the second stage after the realization of the uncertain variables, and they are referred as the “wait-and-see” decisions. Thirdly, $u$ stands for the uncertainty variables which are the outputs of the wind turbines and the PVs, i.e. $P_{\text{WT}}$ and $P_{\text{PV}}$, and the uncertain load demand $P_{D}$.
@@
The unified objective (IV-A) is reformulated in the TSRO form as (IV-C1).
@@
$+{C_{\text{Grid}}}-{C_{rev}}+\lambda\cdot\text{Regret}_{\text{SPO}}$ (37)
In the first stage, the variables $\alpha_{ch}$} and $\alpha_{dis}$} for energy storage are optimized to minimize the cost $C_{\text{ESS}}$. The max operation then considers the worst-case realizations of $P_{\text{WT}}$ and $P_{\text{PV}}$ from these uncertainty sets, capturing adversarial scenarios. Finally, the inner min operation adjusts operational variables $K_{\text{DLC}}$, $V$, $P$ and $Q$ to minimize the total cost, including $C_{\text{WT}}+C_{\text{PV}}$, grid exchange costs $C_{\text{grid}}-C_{\text{rev}}$ and a regret term $\lambda\cdot\text{Regret}_{\text{SPO}}$.
@@ IV-C2 Column-and-Constraint Generation (C&CG) Algorithm
To address these issues, we propose an augmented C&CG method that incorporates a bi-level optimization structure which retains the decomposition for computational tractability while allowing for SPO gradient updates. Additionally, it includes an online learning step within the C&CG loop to update the forecasting model parameters using the SPO gradient derived from the operational decisions.
@@
The SP aims to identify the worst-case uncertainty realization for the given first-stage solution $x^{*}$, which yields the worst-case scenario $u^{*}$ and establishes an upper bound for the objective value. The SP essentially evaluates the robustness of $x^{*}$ by stress-testing it against the most adverse uncertainty realization within $\mathcal{U}$.
@@
To slove the SP, this paper utilize the Karush-Kuhn-Tucker (KKT) optimality conditions and Big-M method to convert the bi-level problem into a MILP model. The reformulation becomes:
$\max_{u,y,\mu,\nu}\;\;$ $d^{T}y+e^{T}u+\lambda\cdot\mathrm{Regret}_{\mathrm{SPO}}(u,x^{*},y)$ (48)
@@
The Big-M method provides an effective approach to linearize the nonlinear complementarity constraints arising from the KKT conditions-based reformulation. This technique introduces binary variables and sufficiently large constants to transform the complementary slackness conditions into equivalent linear constraints.
For each complementarity constraint in the form:
$\mu_{i}\left(v_{i}-F_{i}x^{*}-G_{i}y-H_{i}u\right)=0$ (55)
The nonlinear constraint can then be replaced by the following set of linear constraints:
@@
IV-C3 End-to-end Training Architecture
The end-to-end training architecture proposed in this study integrates the MED and the TSRO in a unified pipeline, enabling joint optimization of forecasting accuracy and operational performance. The framework establishes a differentiable pathway between the two models, allowing gradients to flow through the entire system during training, and ensuring that both forecasting and optimization tasks are aligned with decision-making objectives.
@@ 21. Initialization:
3 Initialize MED network parameters $\theta^{(0)}$;
4 Set initial uncertainty sets $\mathcal{U}^{(0)}_{WT},\mathcal{U}^{(0)}_{PV},\mathcal{U}^{(0)}_{D}$ based on nominal forecasts;
5 Set iteration counter $k=0$;
6 2. Forward Propagation: Forecast $\rightarrow$ Optimization $\rightarrow$ Decision Evaluation
@@ 5 Set iteration counter $k=0$;
12Step 3: Solve TSRO via C&CG:
@@ 14    $\min_{x,\eta}\;c^{\top}x+\eta$
16    $\max_{u}\min_{y}\left(d^{\top}y+e^{\top}u+\lambda\,\text{Regret}_{\text{SPO}}(u,x,y)\right)$
@@
18[0.2em]   Obtain optimal cost $J^{*}(\hat{\mathbf{Y}}^{(k)})$ and decisions $x^{*},y^{*}$;
@@
21 3. Backward Propagation: Decision Gradient $\rightarrow$ Forecast Model
22
23Step 5: Compute SPO gradient $\nabla_{\theta}\mathcal{L}_{SPO}$
24[$\left[-\nabla^{2}J^{*}(\hat{\mathbf{Y}}^{(k)})(\mathbf{Y}-\hat{\mathbf{Y}}^{(k)})\right]\cdot\nabla_{\theta}\hat{\mathbf{Y}}^{(k)}$;
25Step 6: Update MED parameters
@@ V-A Dataset Description and Environment Setting
The dataset includes multivariate time-series data collected from urban microgrid networks with a temporal resolution of 15-min intervals. Key features are categorized as follows:
@@ V-B Probabilistic Forecasting Performance
As shown in Fig. 5, the MED model generates probabilistic forecasts for three variables, including PV power, WT power, and load demand, over a 5-minute resolution timeline. The plots exhibit the predicted probability distributions through shaded confidence intervals (e.g., 90% prediction intervals) and point forecasts (median values), compared against actual observed data points. Notably, the MED model demonstrates high accuracy in tracking temporal patterns, with narrow prediction intervals indicating reduced epistemic uncertainty, particularly during peak demand periods and intermittent renewable generation phases. This performance is attributed to the hierarchical architecture of MED, which leverages residual blocks and encoder-decoder chains to fuse historical data with external features. Moreover, the incorporation of the SPO loss function ensures that forecasting errors are minimized in a decision-aware manner, leading to improved alignment with downstream operation.
@@ V-C2 Robustness Evaluation
Fig. 7 presents the variation of total expected operation cost of 69-bus microgrid with the robust uncertainty budget $\Gamma$ for all forecasting–optimization schemes, with shaded bands indicating the 95% confidence intervals derived from 500 Monte Carlo simulations of forecast uncertainty realizations. The uncertainty budget jointly represents the variability in load demand, photovoltaic, and wind power forecasts. The highlighted point at $\Gamma=0.1$, which represents a typical uncertainty level. As $\Gamma$ increases, all methods exhibit higher operating costs, but the degree of increase and the width of confidence intervals differ markedly. The baseline Forecast + TSRO shows the steepest growth and the largest variability, indicating higher sensitivity to uncertainty, whereas the proposed MED + SPO + TSRO maintains both the lowest mean cost, and the 95% CI width reduced by 42% compared with baseline.

### v2

@@ III-B1 Statistical Error Metric
Note that the sample-based CRPS in (9) involves absolute-value terms; in our implementation we use automatic differentiation with subgradients at the non-differentiable points, which does not affect gradient-based training in practice.
III-B2 Surrogate Decision-focused Loss and Differentiability
Traditional probabilistic metrics such as CRPS only measure statistical accuracy, while microgrid operation performance is determined by the downstream optimization decisions. Let $D$ denote the realized uncertain quantity, and $\hat{D}$ denote the forecast output. We write the downstream operation problem in a compact parametric form as
$J(D)=\min_{z\in\mathcal{Z}(D)}f(z;D),\quad z=(x,y)$ (10)
where $x$ and $y$ represent the first and second stage operation variables, respectively, and $\mathcal{Z}(D)$ denotes the feasible set induced by network, ESS and DLC constraints.
@@ III-B2 Surrogate Decision-focused Loss and Differentiability
To obtain a well-defined gradient pathway, we introduce a continuous convex surrogate of the operation layer by relaxing binary variables to $[0,1]$ and applying standard convex relaxations to nonconvex constraints, yielding a convex feasible set $\bar{\mathcal{Z}}(D)$, and adding a Tikhonov regularization term with $\rho>0$ to ensure strong convexity:
$\bar{z}_{\rho}(D)=\arg\min_{z\in\bar{\mathcal{Z}}(D)}f(z;D)+\frac{\rho}{2}\|z\|_{2}^{2}$ (11)
$\bar{J}_{\rho}(D)=f(\bar{z}_{\rho}(D);D)+\frac{\rho}{2}\|\bar{z}_{\rho}(D)\|_{2}^{2}$ (12)
This surrogate is always solvable under mild feasibility and boundedness conditions, and admits a unique optimizer due to $\rho$-strong convexity.
We define a regret-type decision-focused loss by comparing the operational performance under the forecast-induced decision against the oracle decision under the realized $D$:
$L_{\mathrm{SPO}}(D,\hat{D})=f\!\left(\bar{z}_{\rho}(\hat{D});\,D\right)-f\!\left(\bar{z}_{\rho}(D);\,D\right)$ (13)
To improve numerical stability and keep the loss nonnegative and smooth, we further use the softplus operator:
@@
[…]
The gradients are backpropagated using (15) only through the continuous surrogate optimizer in (11)–(12). The TSRO solved by C&CG is used for forward decision evaluation and benchmarking, which is not differentiated.
@@ IV Robust Operation Model
This section introduces the robust operation model used to optimize microgrid performance under uncertainty. The model leverages TSRO to account for uncertainty in the system’s components. Additionally, this section presents the operation constraints, including energy storage and demand-side management, and discusses the solution approach through the Column-and-Constraint Generation (C&CG) algorithm for improved computational efficiency.
@@ IV-A Objective Function
Based on the model established above, the operation objective function is formulated as follows,
$\text{min }\;C_{\mathrm{ESS}}+$ $C_{\mathrm{WT}}+C_{\mathrm{PV}}+C_{\mathrm{Grid}}-C_{\mathrm{rev}}$ (16)
@@
The decision-focused term is used only for training the forecasting model. Specifically, the training loss is
$\min_{\theta}\;\mathbb{E}\Big[\mathrm{CRPS}(\theta)\;+\;\lambda\cdot\mathrm{Regret}_{\mathrm{SPO}}(D,\hat{D}_{\theta})\Big]$ (22)
where $\hat{D}_{\theta}$ is the forecast output of MED with parameters $\theta$, and $\lambda$ is selected by grid search.
@@ IV-C1 TSRO Model
The proposed microgrid coordination model can be converted to a TSRO model, which adopts a two-stage “min–max–min” structure, where first-stage variables $x$ represent pre-dispatch decisions, second-stage variables $y$ capture recourse actions, and $u$ denotes uncertain renewable and load realizations:
@@
Assume that (i) the uncertainty set $\mathcal{U}$ is nonempty, compact and polyhedral; (ii) the first-stage feasible set $\mathcal{X}$ is nonempty and compact; and (iii) the second-stage recourse set is nonempty and bounded for every $(x,u)\in\mathcal{X}\times\mathcal{U}$.
Under the above assumption, the TSRO problem admits an optimal solution and has a finite optimal value. Moreover, the differentiable surrogate problem in (11)-(12) admits a unique optimizer for any input due to $\rho>0$.
Therefore, we classify the three variables in the two stages as $x$, $y$ and $u$. First, $x$ represents a set of decision variables in the first stage operation, including ESS daily charging and discharging binary variables $\alpha_{\text{ch}}$}, $\alpha_{\text{dis}}$}, and DLC contracts for the entire scheduling period $K_{D,\text{con}}$}. Secondly, $y$ stands for a group of adjustable variables in the second stage including the quarter-hour ahead DLC controllable variables. $K_{\text{DLC}}$} for each quarter hour can be optimized in the second stage after the realization of the uncertain variables, and they are referred as the “wait-and-see” decisions. Thirdly, $u$ stands for the uncertainty variables which are the outputs of the wind turbines and the PVs, and the uncertain load demand $P_{D}$.
@@
The unified objective (16) is reformulated in the TSRO form as (IV-C1).
@@
$+{C_{\text{Grid}}}-{C_{rev}}$ (41)
In the first stage, the variables $\alpha_{ch}$} and $\alpha_{dis}$} for energy storage are optimized to minimize the cost $C_{\text{ESS}}$. The max operation then considers the worst-case realizations of $P_{\text{WT}}$ and $P_{\text{PV}}$ from these uncertainty sets, capturing adversarial scenarios. Finally, the inner min operation adjusts operational variables $K_{\text{DLC}}$, $V$, $P$ and $Q$ to minimize the total cost, including $C_{\text{WT}}+C_{\text{PV}}$ and grid exchange costs $C_{\text{grid}}-C_{\text{rev}}$.
@@ IV-C2 Column-and-Constraint Generation (C&CG) Algorithm
To address these issues, we solve TSRO using the standard C&CG procedure, and update the forecasting model using the surrogate SPO gradient.
@@
The SP aims to identify the worst-case uncertainty realization for the given first-stage solution $x^{*}$:
@@
To slove the SP, this paper utilize the KKT optimality conditions and Big-M method to convert the bi-level problem into a MILP model. The reformulation becomes:
$\max_{u,y,\mu,\nu}\;\;$ $d^{T}y+e^{T}u$ (50)
@@
The Big-M method provides an effective approach to linearize the nonlinear complementarity constraints arising from the KKT conditions-based reformulation.
For each complementarity constraint in (56), the nonlinear constraint can then be replaced by the following set of linear constraints:
@@ 21. Initialization:
3 MED network parameters $\theta^{(0)}$;
4 Uncertainty sets $\mathcal{U}^{(0)}_{WT},\mathcal{U}^{(0)}_{PV},\mathcal{U}^{(0)}_{D}$;
5 Iteration counter $k=0$;
6 2. Forward Propagation:
@@ 5 Iteration counter $k=0$;
12Step 3: Solve TSRO via C&CG (forward pass only):
@@ 14    $\min_{x,\eta}\;c^{\top}x+\eta$
16    $\max_{u}\min_{y}\left(d^{\top}y+e^{\top}u\right)$
@@
18[0.2em]   Obtain optimal cost $J^{*}(\hat{\mathbf{Y}}^{(k)})$ and decisions;
@@
[…]
22 3. Backward Propagation:
23
24Step 5: Obtain $\nabla_{\hat{Y}}\mathrm{Regret}_{\mathrm{SPO}}^{(k)}$ (surrogate layer only):
[…]
26Step 6: Update MED parameters:
@@
IV-C3 End-to-end Training Architecture
The end-to-end training architecture proposed in this study integrates the MED and the TSRO in a unified pipeline, enabling joint optimization of forecasting accuracy and operational performance. The framework establishes a differentiable pathway between the two models, allowing gradients to flow through the surrogate decision layer during training, and ensuring that both forecasting and optimization tasks are aligned with decision-making objectives.
@@ V-A Dataset Description and Environment Setting
The dataset includes multivariate time-series data collected from urban microgrid networks with a temporal resolution of 15-min intervals [medtsro]. Key features are categorized as follows:
@@ V-B Probabilistic Forecasting Performance
As shown in Fig. 5, the MED model generates probabilistic forecasts for three variables, including PV power, WT power, and load demand, over a 5-minute resolution timeline. The plots exhibit the predicted probability distributions through shaded confidence intervals and point forecasts (median values), compared against actual observed data points. Notably, the MED model demonstrates high accuracy in tracking temporal patterns, with narrow prediction intervals indicating reduced epistemic uncertainty, particularly during peak demand periods and intermittent renewable generation phases. This performance is attributed to the hierarchical architecture of MED, which leverages residual blocks and encoder-decoder chains to fuse historical data with external features. Moreover, the incorporation of the SPO loss function ensures that forecasting errors are minimized in a decision-aware manner, leading to improved alignment with downstream operation.
@@ V-C2 Robustness Evaluation
Fig. 7 presents the variation of total expected operation cost of 69-bus microgrid with the robust uncertainty budget $\Gamma$ for all forecasting–optimization schemes, with shaded bands indicating the 95% confidence intervals derived from Monte Carlo simulations of forecast uncertainty realizations. The uncertainty budget jointly represents the variability in load demand, photovoltaic, and wind power forecasts. The highlighted point at $\Gamma=0.1$, which represents a typical uncertainty level. As $\Gamma$ increases, all methods exhibit higher operating costs, but the degree of increase and the width of confidence intervals differ markedly. The baseline Forecast + TSRO shows the steepest growth and the largest variability, indicating higher sensitivity to uncertainty, whereas the proposed MED + SPO + TSRO maintains both the lowest mean cost, and the 95% CI width reduced by 42% compared with baseline.